### v2

# references.bib

Abstract
Computational phenotyping is essential for biomedical research but often requires significant time and resources, especially since traditional methods typically involve extensive manual data review. While machine learning and natural language processing advancements have helped, further improvements are needed. Few studies have explored using Large Language Models (LLMs)for these tasks despite known advantages of LLMs for text-based tasks. To facilitate further research in this area, we developed an evaluation framework, Evaluation of PHEnotyping for Observational Health Data (PHEONA), that outlines context-specific considerations. We applied and demonstrated PHEONA on concept classification, a specific task within a broader phenotyping process for Acute Respiratory Failure (ARF)respiratory support therapies. From the sample concepts tested, we achieved high classification accuracy, suggesting the potential for LLM-based methods to improve computational phenotyping processes.
Introduction
Clinical phenotyping, the process of extracting information from health data to classify patients based on relevant characteristics, is fundamental for biomedical research[hripcsak_next-generation_2013]. One such type of these processes is electronic or computational phenotyping, which involves defining cohorts algorithmically using data from electronic health records (EHRs)and other relevant information systems[callahan_characterizing_2023]. The resulting phenotypes have been used for many clinical outcomes and conditions supporting a variety of downstream tasks, including recruitment for clinical trials, development of clinical decision support systems, and hospital quality reporting[banda_advances_2018, tekumalla_towards_2024, shang_making_2019]. The process of developing computable phenotypes includes constructing relevant data elements for classification and then applying an algorithm to produce the cohort(s) of interest[carrell_general_2024]. For example, a computable phenotype for Post-Acute Sequelae of SARS-CoV-2 infection (PASC)was developed by first identifying standard clinical concepts for each PASC-related symptom and then applying an algorithm to identify individuals with PASC based on symptom presence relative to SARS-CoV-2 infection using EHR data[pungitore_computable_2024].
Traditionally, development of computable phenotypes requires time-intensive, manual data review, including identification of relevant data and potential mapping to controlled vocabularies[shang_making_2019]. From the PASC example, 6,56965696,5696 , 569 concepts across $151151$ symptoms were manually reviewed over several weeks by a clinician expert to determine relevance of each concept to PASC-specific symptoms[pungitore_computable_2024]. While developments in Natural Language Processing (NLP)and machine learning methods could improve processing and categorization of clinical data elements, these methods also add additional steps such as model training and hyperparameter tuning which can be similarly time and resource-intensive[banda_advances_2018, shang_making_2019]. Thus, there is a need for substantial improvements to traditional computational phenotyping processes, particularly for tasks requiring manual review. As a relatively new addition to biomedical research, Large Language Models (LLMs)introduce a brand new set of text analysis and generation capabilities that could supplement or replace manual review of clinical data elements[raiaan_review_2024]. Furthermore, LLMs can be trained on new tasks through prompt engineering alone, which is a more accessible and portable method of model adaptation compared to traditional model retraining or domain adaptation methods[liu_pre-train_2023]. Therefore, we hypothesize LLMs can transform computational phenotyping, in particular through acceleration and enrichment of manual review through incorporation of LLM-supported processes.
To our knowledge, there have been no studies that have explored the application of LLMs to computational phenotyping tasks. Therefore, this study makes the following contributions:
  1. 1.
    We developed PHEONA (Evaluation of PHEnotyping for Observational Health Data), a framework that outlines an evaluation mechanism for LLM-based computational phenotyping.
  2. 2.
    We provided an example usage of all components of PHEONA to promote further research and advancements in computational phenotyping processes.
  3. 3.
    We demonstrated a successful application of LLM-based methods to concept classification, a step within a broader phenotyping process that was previously performed solely via manual review.
Methods
We first discuss how individual components of PHEONA (Figures 1 and 2) were constructed, including descriptions of the associated criteria. We then outline how we demonstrated PHEONA on a specific phenotyping use case.
Figure 1: Overview of the components of Evaluation of PHEnotyping for Observational Health Data (PHEONA).
Development of PHEONA
Due to the novelty of applying LLMs to computational phenotyping, the first component of PHEONA, the resource requirements, was designed to determine when LLM-based methods may provide advantages over traditional phenotyping methods. The resource requirements outline considerations to estimate time, cost, and computational effort and should be considered before development begins. We based the resource requirements on components of Constructive Cost Model (COCOMO)$2.02.0$, a cost estimation model for software development projects[boehm_cost_1995]. Similar to COCOMO, we focused on attributes of the development effort (amount of work required) and development schedule (timeline of work required) when considering resource requirements for LLM-based projects. However, unlike COCOMO, we assumed an experienced research team so that development effort is primarily a function of existing hardware and software constraints. Table 1 outlines individual attributes of the development effort and schedule for consideration.
Following an assessment of whether LLMs could be considered for a specific task, the second component of PHEONA, model ability, provides criteria to evaluate LLM-specific processes (i.e., prompt engineering) and known drawbacks of LLMs (i.e., hallucination and response instability)[li_llms-as-judges_2024, zhuo_prosa_2024, patil_review_2024]. Therefore, model ability includes the following subcategories: 1) Prompt Factors; 2) Model Consistency; and 3) Response Correctness. Prompt factors evaluate the effort required for prompt engineering as well as the corresponding response times for each prompt. Model consistency factors evaluate the ability of the model to provide consistent answers, even if changes are made to the prompt, since LLMs can be sensitive to inputs[li_llms-as-judges_2024, zhuo_prosa_2024]. Additionally, automated response extraction from LLM responses is necessary for large volume requests and responses must be consistent to produce reliable results. Finally, response correctness factors were designed to evaluate overall performance of the model, including the presence of hallucinations. Table 2 details the model ability requirements. We note that when evaluating the model ability criteria, a random subset of the entire dataset may be necessary for evaluation, especially for large datasets without previously defined ground truths.
Phenotyping Use Case
We applied PHEONA to concept classification, a single development step in a previously defined computable phenotype for Acute Respiratory Failure (ARF)respiratory support therapies. Although the phenotypes were discussed in-depth previously, we provide a brief overview to contextualize our work[essay_rule-based_2020]. The phenotyping algorithm classified individuals based on the respiratory therapies received, specifically Invasive Mechanical Ventilation (IMV), High-Flow Nasal Insufflation (HFNI), and Noninvasive Positive Pressure Ventilation (NIPPV), during an individual Intensive Care Unit (ICU)stay. The respiratory therapies were identified by the presence of specific concepts, such as “Hi Flow NC” and “BiPAP/CPAP”, that indicated therapy use. Additionally, for IMV, presence of sedation and paralysis medications was required to verify patients underwent intubation. In the original study, manual review of the data elements was performed to determine which concepts were relevant to any of the respiratory therapies or specific medications. For our LLM-based classification, we engineered prompts to determined which concepts were relevant to at least one of the following categories: IMV, HFNI, NIPPV, and medications used for sedation and paralysis.
Figure 2: Overview of how to apply Evaluation of PHEnotyping for Observational Health Data (PHEONA)to a computational phenotyping task.
Data Processing
We used the eICU Collaborative Research Database (eICU-CRD)dataset since concept ground truths existed prior to this study from manual review of the eICU-CRD data during development of the ARF respiratory support therapy phenotypes[essay_rule-based_2020, pollard_eicu_2018]. EHR data were processed to generate constructed data concepts using eICU-CRD tables with timestamped patient data that likely contained information on respiratory therapies based on the table descriptions. Tables that did not include timestamped data but contained information on respiratory therapies (such as the apacheApsVar table) or that were unlikely to contain descriptions of respiratory therapies (such as the vitalPeriodic table) were excluded from further processing. There were $99$ total tables selected that were used to construct input concepts. Each constructed concept followed the format: Source: “Table Name”; Concept = “Concept Name and Details”, where Table Name was the name of the input table (e.g., Infusion Drug), and Concept Name and Details was a string concatenation of potentially relevant columns (see Table 3).
Prompt Engineering
Prompts were developed and refined using a random sample of $100100$ constructed concepts. Since best results on the sample were obtained when categorization of respiratory therapies (IMV, NIPPV, and HFNI) was separate from categorization of sedation and paralysis medications, we developed two prompts, resulting in two LLM responses per constructed concept[pungitore_prompts_2025]. The prompts were structured into multiple sections, including input, instructions, concept definitions, and output. Each constructed concept was injected into the input section. After prompt development, we required only the output section of the prompt for model ability evaluation, which included a series of questions and answers to guide the model into determining whether the constructed concept was relevant to either the respiratory therapies or medications of interest. The answer to the final question for each prompt was parsed using string methods to get the final response of “YES” or “NO” for whether the constructed concept was relevant. Since there were two prompts used, the final response was “YES” if at least one of the responses was “YES” and “NO” otherwise.
Model Setup
The models tested were identified from popular models available through Ollama, an open-source package that allows hosting of various open-source LLM models locally[noauthor_ollamaollama_2024]. Due to graphical processing unit (GPU)resource constraints, we selected the following instruction-tuned models for analysis: Mistral Small $2424$ billion with Q8.08.08.08.0 quantization (model tag: 20ffe5db0161), Phi-$44$ $1414$ billion with Q8.08.08.08.0 quantization (model tag: 310d366232f4), Gemma2222 $2727$ billion with Q8.08.08.08.0 quantization (model tag: dab5dca674db), and DeepSeek-r1 $3232$ billion with Q4444_K_M quantization (model tag: 38056bbcbb2d)[noauthor_ollamaollama_2024]. Models were run on either a single Nvidia V100100100100 32323232GB or two Nvidia P100100100100 16161616GB GPUs depending on server availability. Temperature and top-p were set at $0.00.0$ and $0.990.99$, respectively, for all experiments.
Assessment of Resource Requirements
The resource requirements were assessed by comparing manual review of the original 83,2898328983,28983 , 289 concepts to automated LLM classification of the same concepts. In the original study, manual review of all concepts was performed by $33$ experts over an estimated $66$ months out of the total $1818$ months required for phenotype development. Development effort and schedule for the LLM-based approach were measured during prompt engineering for the initial $100100$ samples.
Assessment of Model Ability
To assess model ability, we first evaluated all components of PHEONA to fully demonstrate the framework and then selected the best model for concept classification based on results specifically for response correctness, response latency, and response format accuracy. We prioritized response correctness to ensure accurate classification results. Response latency and response format accuracy were secondary evaluation metrics to ensure the selected model could quickly and consistently categorize the entire set of 83,2898328983,28983 , 289 concepts. The time to minimal viable prompt was measured on the initial sample of $100100$ constructed concepts for prompt engineering while the remaining metrics were measured using another random sample of $100100$ constructed concepts to avoid data leakage. We demonstrated the model ability criteria using data samples to mimic phenotype development where ground truths are unlikely to be available and it would be necessary to manually classify concepts prior to response correctness calculations.
For prompt factors, the time to minimum viable prompt was measured as the time (in hours) for prompt engineering of the prompts for classification of the sample constructed concepts. Model response latency was calculated by first summing the response latencies across the two API calls required for each concept and then averaging the latency across the $100100$ samples. Since we used internal environments for hosting the LLMs, we did not consider testing at different times of the day. For model consistency, the response format accuracy was determined by counting the number of responses from both prompts that were returned in the exact output format specified in the prompts[pungitore_prompts_2025]. We then tested the response consistency and prompt stability by first randomly selecting $1010$ constructed concepts from the original sample of $100100$. For response consistency, we generated $1010$ model results per constructed concept and evaluated how many times the model returned the same answer (either “YES” or “NO”) for each one for a total of $200200$ responses per model. The prompt stability was tested by rearranging the respiratory therapy categorization prompt (since it was large and complex) and evaluating whether the model return the same answer for each constructed concept. The changes made to the prompt were the following: 1) Moved the instructions after the concept criteria; 2) Reversed the ordering of the concept-related questions; and 3) Reversed the ordering of the concepts within the prompt. Each change was independently tested $1010$ times on each of the $1010$ sampled constructed concepts for a total of $200200$ responses per change.
For response correctness, the accuracy was measured using the area under the receiver operating characteristic (AUROC)score. The quasi-accuracy was also measured with AUROC after reclassifying any relevant false positives as true positives based on the prompt concept definitions after manual review[pungitore_prompts_2025]. Hallucination frequency was determined by assessing the number of factually incorrect statements, such as incorrect acronym interpretations or incorrect association of medications to their functional group. The severity of each hallucination was evaluated as follows: Minor (little to no impact on the model reasoning or final decision); Major (impact on model reasoning but not the final decision, i.e., the right decision made for the wrong reason); Critical (impact on both model reasoning and final decision).
Results
Assessment of Resource Requirements
The resource requirements comparison for both methods is presented in Table 4. From this comparison, we note that while the LLM-based methods required more specialized hardware and software, the time savings for classification was estimated to be significant when compared to manual review in the traditional approach, especially when considering that the actual concept classification is performed automatically by the LLM. Therefore, we decided LLM-based methods would be appropriate to explore for concept classification.
Assessment of Model Ability
For time to minimum viable prompt, early testing showed that a prompt developed on one model was effective for the remaining models and therefore, we did not estimate any differences in this metric between models. For the response correctness metrics, there were no hallucinations detected across any of the models. Additionally, all of the models achieved an AUROC of $1.0001.000$ for quasi-accuracy after reviewing the concepts identified as false positives. The false positives reassigned as true positives were Source = Nurse Charting; Concept = O2 Admin Device: BiPAP/CPAP (all models) and Source = Nurse Charting; Concept = O2 Admin Device: nasal cannula (Gemma and Mistral, only) since they were both relevant to the respiratory therapy concepts of NIPPV and HFNI, respectively, as defined in the classification prompts[pungitore_prompts_2025]. The results for the remaining metrics are presented in Table 5. DeepSeek had the highest model response latency ($40.040.0$ seconds compared to $17.217.2$ seconds or faster) and lowest response format accuracy (149/200149200149/200149 / 200 compared to 177/200177200177/200177 / 200 or better). Additionally, response format accuracy was lower for Gemma (177/200177200177/200177 / 200) when compared to Mistral (197/200197200197/200197 / 200) or Phi (194/200194200194/200194 / 200), indicating Gemma was less suitable for concept classification. Since Mistral and Phi had similar metrics for model response latency and response format accuracy, we would select either of these models for categorization of the entire constructed concept dataset.
Discussion
In this study, we developed PHEONA, an evaluation framework specifically for LLM applications to computational phenotyping. In addition to outlining components of PHEONA, we demonstrated how to apply it to concept classification, a single task within a broader phenotyping pipeline for ARF respiratory support therapies.
Evaluation Framework
The primary contribution of this study was development of PHEONA to support advancement of computational phenotyping methods by integrating LLMs. Despite recent improvements, computational phenotyping remains complicated and costly[shang_making_2019, tekumalla_towards_2024] and many of the desiderata for computational phenotyping identified over a decade ago are still relevant today, indicating the continued need for substantial improvements[hripcsak_next-generation_2013, wen_impact_2023]. Foundation models like LLMs are the first widely available tools with the capability to comprehend and produce human-readable text for a fraction of the development time and thus, they are a natural next step for computational phenotyping. With PHEONA, we provided guidelines to support integration of LLMs into phenotyping pipelines. The resource requirements component of PHEONA was designed to assess whether LLMs are appropriate for the task of interest since not every task could or should use an LLM based on available resources. The model ability component was then designed to determine if the available LLMs could accomplish the task of interest. To simplify development and avoid additional costs, we recommend researchers start with smaller parameter LLMs and consider larger models if performance is low. There are other methods to consider, such as fine-tuning or multiple model approaches[li_llms-as-judges_2024], but these may not provide an advantage over traditional methods and should be evaluated using resource requirements criteria on a case-by-case basis. In scenarios where the resource requirements are revisited, the application of PHEONA would look similar to Figure 2, where the resource requirements and model ability are cycled through to find the optimal method for a given task.
Phenotyping Use Case
Another contribution of this study was the initial results of concept classification for ARF respiratory support therapies. From our evaluation of the model ability criteria, we concluded that we could accomplish concept classification effectively with either Mistral or Phi, although the models we tested performed well at classifying the constructed concept sample. All of the tested models had high response correctness scores and thus, we only focused on secondary metrics (model response latency and response format accuracy) for final model evaluation. With respect to these metrics, DeepSeek performed marginally worse than the other models; however, it was also the only model to incorporate reasoning into the response generation, which makes the responses less consistent and more lengthy but potentially more valuable for complex tasks. These results demonstrated several advantages of LLM-based approaches specifically for classification tasks. First, these results showed that high accuracy can be achieved without extensive feature engineering or model training. Second, our methods incurred no additional hardware costs, indicating that LLM-based methods do not need expensive workflows to be effective for classification. Finally, our results were achieved with minimal data processing and relatively fast latency, suggesting potential further adaptation of LLMs to time-sensitive or data-intensive clinical workflows.
Study Limitations
There are some limitations to the framework and methods identified in this study. First, there are known issues with using LLMs, including the ability to produce hallucinations and reduced performance around specialized, domain-specific knowledge[li_llms-as-judges_2024, patil_review_2024]. While the model ability components of PHEONA help highlight where LLMs may fail, there may be cases where LLMs cannot accomplish the task of interest and traditional phenotyping methods must be used instead. Additionally, the current version of PHEONA requires manual review of responses to measure each criterion, despite the known drawback of manual review being time-intensive. Although we used a random sample of the total constructed concepts to make manual review more feasible, it is possible that performance on the entire dataset is not representative of performance on the random sample. Finally, for our use case, we only considered the binary classification task of “YES/NO” for whether each constructed concept was relevant to the topics of interest although multi-class classification may be more useful for downstream processes. However, when reviewing the responses to individual classification steps in our prompt, we did not find any inaccurate responses and thus, we should be able to convert from binary to multi-class classification with some minor prompt updates.
Future Directions
There were several future directions identified from this study. First, since one of the limitations of this study was the reliance on manual review for evaluation, there are opportunities to develop automated LLM processes for computational phenotyping and to update PHEONA to incorporate automated evaluation processes. Furthermore, while we demonstrated PHEONA on a single task, the framework can be generalized to adapt LLM-based methods to entire computational phenotyping pipelines rather than simply adapting LLMs to individual phenotyping tasks. Finally, each of the criteria in PHEONA can be expanded upon to include a more standardized evaluation approach. For example, there have been frameworks developed to measure LLM sensitivity to general inputs that can be adapted to the prompt stability component of PHEONA [zhuo_prosa_2024].
Conclusions
We developed PHEONA, an evaluation framework for applying LLM-based methods to computational phenotyping tasks. This framework can be used to assess whether LLMs are feasible and if so, which model is best for the task of interest. We provided an example use case to demonstrate the framework and promote further research into improving traditional computational phenotyping methods.
[…]
Table 1: Evaluation criteria for resource requirements, specifically attributes of development effort and development schedule. These criteria were designed to compare the necessary resources for LLM-based approaches to traditional phenotyping approaches.
Development Effort
Criterion Description How to Measure
Model Host Environments Availability of external and internal environments for hosting models and artifacts based on project budget, expected timeline, Protected Health Information (PHI)restrictions, and/or available compute environments. Consider each restriction from most to least significant and use to determine the list of feasible environments to host models. For example, use of PHI may limit access to externally hosted models if privacy safeguards cannot be guaranteed through standard end-user license agreements.
Hardware Requirements Minimum hardware requirements required to run model(s) of interest if considering internal environments. May be re-evaluated to accommodate any inference optimization strategies implemented[chitty-venkata_survey_2023]. Assess the maximum computational resources in each of the available host environments, focusing on graphical processing unit (GPU)resources in particular. Assess whether maximum resources will be available throughout the project or if a smaller model size is desired to ensure consistent access.
Software Availability For externally hosted models, consideration of the availability of software to interact with these models. For internally hosted models, consideration of the availability and complexity of software to host and interact with selected models. Assess the software requirements to interact with models in each of the available host environments and determine time to setup and develop with available software.
Development Schedule
Criterion Description How to Measure
Time for Pipeline Development The amount of time required to implement the desired phenotyping methods, including development and evaluation time. Estimate the required development time to complete each of the phenotyping steps. For LLM-based methods these steps would likely encompass model setup, data processing, prompt engineering, and response evaluation.
Time for Manual Review The amount of time required for manual review of responses based on the phenotyping task and overall pipeline. Based on the steps in the phenotyping pipeline that would require manual review, estimate total review time based on quantity of data and any initial results.
Phenotype Runtime The amount of time required to perform all included phenotyping tasks on one record, such as data processing time, model runtime, and/or any application programming interface (API)response times. Estimate runtimes based on size of data and complexity of phenotyping process. For model API calls, use example data and prompts to estimate total phenotyping time. If using an externally hosted model or other external resources, consider testing APIs at different times of the day to test latency variations.
Table 2: Evaluation criteria for each subcategory under model ability. These criteria were designed to determine the model that achieves optimal results on specific phenotyping tasks.
Prompt Factors
Criterion Description How to Measure
Time to Minimum Viable Prompt The amount of time to engineer prompt with acceptable results. Estimate prompt engineering time after testing different prompts and prompt structures (such as few-shot learning [liu_pre-train_2023]), considering how adaptable each model is to increasing prompt complexity.
Model Response Latency The time to return responses for each prompt. Measure the average amount of time to receive responses from the LLM for each prompt of interest. If using an externally hosted model, consider testing at different times of the day to assess latency variations.
Model Consistency
Criterion Description How to Measure
Response Format Accuracy The accuracy of correctly formatted responses based on the requested format in the prompt. From the requested structure in the prompt, determine how many responses were returned in this format with no deviations.
Response Consistency The consistency of responses with multiple model runs. Run each model multiple times with the same prompt on the same data and record how frequently model returns same answer.
Prompt Stability The ability of the prompts to return consistent responses even if there are perturbations in the prompts. Make changes to the prompt and re-evaluate relevant criteria, such as response format accuracy and classification accuracy. Example perturbations include reordering prompt sections and removing any text emphasis.
Response Correctness
Criterion Description How to Measure
Accuracy How accurate the model is at classifying or phenotyping based on a ground truth. Classification accuracy can be measured using common classification metrics like accuracy, precision, recall, and area under the receiver operating characteristic (AUROC).
Quasi-Accuracy Classification accuracy after marking relevant or semantically similar false positives as true positives to avoid penalizing LLM when unnecessary. Review responses classified as False Positive and determine if either a) the original text is semantically relevant to the classification task of interest or b) there is evidence that the ground truth is incorrect. Recalculate classification metrics after reassigning these cases to True Positives.
Hallucination Frequency and Severity Determine the frequency and corresponding severity of hallucinations in LLM responses. Assess frequency of factually incorrect statements in responses. For each hallucination, assess severity based on, for example, whether the outcome was affected, severe biases were introduced, or the response detracted from the overall goal of the prompt.
Table 3: Constructed concept pattern for each selected table in the eICU Collaborative Research Database (eICU-CRD)database[pollard_eicu_2018]. Italicized text was replaced with specific values from the relevant column in each table.
Table Name Constructed Concept Pattern
Care Plan General Source = Care Plan General; Concept = cplgroup: cplitemvalue
Infusion Drug Source = Infusion Drug; Concept = drugname
Medication Source = Medication; Concept = drugname
Note Source = Note; Concept = notevalue: notetext
Nurse Care Source = Nurse Care; Concept = cellattributevalue
Nurse Charting Source = Nurse Charting; Concept = nursingchartcelltypevalname: nursingchartvalue
Respiratory Care Source = Respiratory Care; Concept = airwaytype
Respiratory Charting Source = Respiratory Charting; Concept = respcharttypecat: respchartvaluelabel: respchartvalue
Treatment Source = Treatment; Concept = treatmentstring
Table 4: Demonstration of the resource requirements to decide between an LLM-based or a traditional phenotyping approach for the task of concept classification in a computable phenotype for Acute Respiratory Failure (ARF)respiratory support therapies.
Development Effort
Criterion Traditional Approach LLM-based Approach Verdict
Model Host Environments Due to PHI concerns, only internal environments were available for use. Due to PHI concerns, only internal environments were available for use. Environments with GPU resources were available but restricted based on demand. Traditional approach was superior due to ability to run model on multiple internal environments rather than being restricted to those with GPU access.
Hardware Requirements No additional computational resources or cost expected. Access to high performance research servers was limited based on demand but incurred no additional cost. Traditional approach was marginally superior due to limits on GPU resources although neither approach incurred additional costs.
Software Requirements No additional software requirements specific to phenotyping required. Software specific to interacting with LLM models was required and troubleshooting was required to adapt software to high performance compute environment. Traditional approach was superior due to extra setup time for LLM software.
Development Schedule
Criterion Traditional Approach LLM-based Approach Verdict
Time for Pipeline Development An estimated $66$ months of the total $1818$ months for ARF phenotype development were required for manual review of concepts. A total of $22$ months was required for prompt engineering on the initial sample of concepts. An estimated additional month would be required to gather all concept responses for all models based on sample latencies with models running in parallel. LLM-based method was superior.
Time for Manual Review An estimated $66$ months of the total $1818$ months for ARF phenotype development were required for manual review of concepts. An week was required for clinical experts to collate information on the concepts of interest to be used for prompt engineering and an additional day was required to manually review model ability results. LLM-based method was superior.
Phenotype Runtime There were no artifacts specifically for concept classification. An estimated $3030$ seconds was required to classify each concept as relevant for the ARF respiratory support phenotypes. Not Applicable, since the traditional phenotyping methods relied only on manual review for concept classification.
Table 5: Demonstration of the model ability criteria to determine best LLM for concept classification. When included, the denominators indicate the number of Large Language Model (LLM)responses that were evaluated for each metric based on the specific experiment implemented.
Prompt Factors
Criterion DeepSeek Gemma Mistral Phi
Model Response Latency $40.040.0$ seconds $17.217.2$ seconds $14.814.8$ seconds $11.311.3$ seconds
Model Consistency
Criterion DeepSeek Gemma Mistral Phi
[…]
Prompt Stability
[…]
Response Correctness
Criterion DeepSeek Gemma Mistral Phi
Accuracy (AUROC a) $0.9950.995$ $0.9900.990$ $0.9900.990$ $0.9950.995$
a AUROC: area under the receiver operating characteristic curve.

### v3

@@
-references.bib
+cleaned_references.bib
 Abstract
@@
 For example, a computable phenotype for Post-Acute Sequelae of SARS-CoV-2 infection (PASC)was developed by first identifying standard clinical concepts for each PASC-related symptom and then applying an algorithm to identify individuals with PASC based on symptom presence relative to SARS-CoV-2 infection using EHR data[pungitore_computable_2024].
-Traditionally, development of computable phenotypes requires time-intensive, manual data review, including identification of relevant data and potential mapping to controlled vocabularies[shang_making_2019].
+Traditionally, development of computable phenotypes requires time-intensive, manual review of data and definitions, including identification of relevant data and mapping to controlled vocabularies[shang_making_2019].
 From the PASC example, 6,56965696{,}5696 , 569 concepts across $151151$ symptoms were manually reviewed over several weeks by a clinician expert to determine relevance of each concept to PASC-specific symptoms[pungitore_computable_2024].
-While developments in Natural Language Processing (NLP)and machine learning methods could improve processing and categorization of clinical data elements, these methods also add additional steps such as model training and hyperparameter tuning which can be similarly time and resource-intensive[banda_advances_2018, shang_making_2019].
-Thus, there is a need for substantial improvements to traditional computational phenotyping processes, particularly for tasks requiring manual review.
+While developments in Natural Language Processing (NLP)and machine learning methods could improve processing and categorization of clinical data, these methods also add additional steps such as model training and hyperparameter tuning which can be similarly time and resource-intensive[banda_advances_2018, shang_making_2019].
+Thus, there is a need for substantial improvements to traditional computational phenotyping processes, particularly for tasks that require manual review or intervention.
 As a relatively new addition to biomedical research, Large Language Models (LLMs)introduce a brand new set of text analysis and generation capabilities that could supplement or replace manual review of clinical data elements[raiaan_review_2024].
@@
 Development of PHEONA
-Due to the novelty of applying LLMs to computational phenotyping, the first component of PHEONA, the resource requirements, was designed to determine when LLM-based methods may provide advantages over traditional phenotyping methods.
+Since LLMs have not previously been applied to computational phenotyping tasks, the first component of PHEONA, the resource requirements, was designed to determine when LLM-based methods may provide advantages over traditional phenotyping ones.
 The resource requirements outline considerations to estimate time, cost, and computational effort and should be considered before development begins.
@@
 Finally, response correctness factors were designed to evaluate overall performance of the model, including the presence of hallucinations.
-Table 2 details the model ability requirements.
+These criteria were most closely related to traditional phenotyping evaluation, where phenotyping performance is measured against ground truths.[carrell_general_2024] Table 2 details the model ability requirements.
 We note that when evaluating the model ability criteria, a random subset of the entire dataset may be necessary for evaluation, especially for large datasets without previously defined ground truths.
 Phenotyping Use Case
-We applied PHEONA to concept classification, a single development step in a previously defined computable phenotype for Acute Respiratory Failure (ARF)respiratory support therapies.
+We applied PHEONA to concept classification, a single development step for previously defined computable phenotypes for Acute Respiratory Failure (ARF)respiratory support therapies.
 Although the phenotypes were discussed in-depth previously, we provide a brief overview to contextualize our work[essay_rule-based_2020].
-The phenotyping algorithm classified individuals based on the respiratory therapies received, specifically Invasive Mechanical Ventilation (IMV), High-Flow Nasal Insufflation (HFNI), and Noninvasive Positive Pressure Ventilation (NIPPV), during an individual Intensive Care Unit (ICU)stay.
+The phenotyping algorithm classified individuals based on the respiratory therapies received, specifically Invasive Mechanical Ventilation (IMV), Noninvasive Positive Pressure Ventilation (NIPPV), and High-Flow Nasal Insufflation (HFNI)during an individual Intensive Care Unit (ICU)stay.
 The respiratory therapies were identified by the presence of specific concepts, such as “Hi Flow NC” and “BiPAP/CPAP”, that indicated therapy use.
@@
 In the original study, manual review of the data elements was performed to determine which concepts were relevant to any of the respiratory therapies or specific medications.
-For our LLM-based classification, we engineered prompts to determined which concepts were relevant to at least one of the following categories: IMV, HFNI, NIPPV, and medications used for sedation and paralysis.
+For our LLM-based classification, we engineered prompts to determine which concepts were relevant to at least one of the following categories: IMV, NIPPV, HFNI, and medications used for sedation and paralysis.
 Figure 2: Overview of how to apply Evaluation of PHEnotyping for Observational Health Data (PHEONA)to a computational phenotyping task.
 Data Processing
-We used the eICU Collaborative Research Database (eICU-CRD)dataset since concept ground truths existed prior to this study from manual review of the eICU-CRD data during development of the ARF respiratory support therapy phenotypes[essay_rule-based_2020, pollard_eicu_2018].
-EHR data were processed to generate constructed data concepts using eICU-CRD tables with timestamped patient data that likely contained information on respiratory therapies based on the table descriptions.
+We used the eICU Collaborative Research Database (eICU-CRD)dataset since concept ground truths existed prior to this study from development of the ARF respiratory support therapy phenotypes[essay_rule-based_2020, pollard_eicu_2018].
+EHR data were processed to generate constructed data concepts using eICU-CRD tables with timestamped data that likely contained information on respiratory therapies based on the table descriptions.
 Tables that did not include timestamped data but contained information on respiratory therapies (such as the apacheApsVar table) or that were unlikely to contain descriptions of respiratory therapies (such as the vitalPeriodic table) were excluded from further processing.
@@
 The answer to the final question for each prompt was parsed using string methods to get the final response of “YES” or “NO” for whether the constructed concept was relevant.
-Since there were two prompts used, the final response was “YES” if at least one of the responses was “YES” and “NO” otherwise.
+Since there were two prompts used, the final response was “YES” if at least one of the responses was “YES” and “NO,” otherwise.
 Model Setup
-The models tested were identified from popular models available through Ollama, an open-source package that allows hosting of various open-source LLM models locally[noauthor_ollamaollama_2024].
-Due to graphical processing unit (GPU)resource constraints, we selected the following instruction-tuned models for analysis: Mistral Small $2424$ billion with Q8.08.08.08.0 quantization (model tag: 20ffe5db0161), Phi-$44$ $1414$ billion with Q8.08.08.08.0 quantization (model tag: 310d366232f4), Gemma2222 $2727$ billion with Q8.08.08.08.0 quantization (model tag: dab5dca674db), and DeepSeek-r1 $3232$ billion with Q4444_K_M quantization (model tag: 38056bbcbb2d)[noauthor_ollamaollama_2024].
+The models tested were identified from popular models available through Ollama, an open-source package that allows local hosting of open-source LLMs [noauthor_ollamaollama_2024].
+Due to graphics processing unit (GPU)resource constraints, we selected the following instruction-tuned models for analysis: Mistral Small $2424$ billion with Q8.08.08.08.0 quantization (model tag: 20ffe5db0161), Phi-$44$ $1414$ billion with Q8.08.08.08.0 quantization (model tag: 310d366232f4), Gemma2222 $2727$ billion with Q8.08.08.08.0 quantization (model tag: dab5dca674db), and DeepSeek-r1 $3232$ billion with Q4444_K_M quantization (model tag: 38056bbcbb2d)[noauthor_ollamaollama_2024].
 Models were run on either a single Nvidia V100100100100 32323232GB or two Nvidia P100100100100 16161616GB GPUs depending on server availability.
@@
 Assessment of Resource Requirements
-The resource requirements were assessed by comparing manual review of the original 83,2898328983,28983 , 289 concepts to automated LLM classification of the same concepts.
+The resource requirements were assessed by comparing manual review of the original 83,2898328983{,}28983 , 289 concepts to automated LLM classification.
 In the original study, manual review of all concepts was performed by $33$ experts over an estimated $66$ months out of the total $1818$ months required for phenotype development.
-Development effort and schedule for the LLM-based approach were measured during prompt engineering for the initial $100100$ samples.
+Development effort and schedule for the LLM-based approach were measured during prompt engineering for the initial sample of $100100$ constructed concepts.
 Assessment of Model Ability
-To assess model ability, we first evaluated all components of PHEONA to fully demonstrate the framework and then selected the best model for concept classification based on results specifically for response correctness, response latency, and response format accuracy.
+To assess model ability, we evaluated all components of PHEONA to fully demonstrate the framework and then selected the best model for concept classification based on response correctness, response latency, and response format accuracy.
 We prioritized response correctness to ensure accurate classification results.
-Response latency and response format accuracy were secondary evaluation metrics to ensure the selected model could quickly and consistently categorize the entire set of 83,2898328983,28983 , 289 concepts.
-The time to minimal viable prompt was measured on the initial sample of $100100$ constructed concepts for prompt engineering while the remaining metrics were measured using another random sample of $100100$ constructed concepts to avoid data leakage.
-We demonstrated the model ability criteria using data samples to mimic phenotype development where ground truths are unlikely to be available and it would be necessary to manually classify concepts prior to response correctness calculations.
-For prompt factors, the time to minimum viable prompt was measured as the time (in hours) for prompt engineering of the prompts for classification of the sample constructed concepts.
-Model response latency was calculated by first summing the response latencies across the two API calls required for each concept and then averaging the latency across the $100100$ samples.
+Response latency and response format accuracy were secondary evaluation metrics to ensure the model could quickly and consistently categorize all concepts.
+The time to minimal viable prompt was measured on the initial sample of $100100$ constructed concepts and the remaining metrics were measured using another random sample of $100100$ constructed concepts to avoid data leakage.
+We demonstrated the model ability criteria using data samples to mimic phenotype development where ground truths are unlikely to be available and it would be necessary to manually classify concepts prior to evaluation.
+For prompt factors, the time to minimum viable prompt was measured as the time to develop prompts for classification of the sample constructed concepts for each model.
+Model response latency was calculated by first summing the response latencies across the two LLM responses and then averaging the latency for the $100100$ samples.
 Since we used internal environments for hosting the LLMs, we did not consider testing at different times of the day.
-For model consistency, the response format accuracy was determined by counting the number of responses from both prompts that were returned in the exact output format specified in the prompts[pungitore_prompts_2025].
-We then tested the response consistency and prompt stability by first randomly selecting $1010$ constructed concepts from the original sample of $100100$.
-For response consistency, we generated $1010$ model results per constructed concept and evaluated how many times the model returned the same answer (either “YES” or “NO”) for each one for a total of $200200$ responses per model.
-The prompt stability was tested by rearranging the respiratory therapy categorization prompt (since it was large and complex) and evaluating whether the model return the same answer for each constructed concept.
+The response format accuracy was determined by counting the number of responses that were returned in the exact output format specified in the prompts[pungitore_prompts_2025].
+We tested response consistency and prompt stability by first randomly selecting $1010$ constructed concepts from the second sample of $100100$.
+For response consistency, we generated $1010$ model results per constructed concept and evaluated how many times the model returned the same answer (either “YES” or “NO”) for a total of $200200$ responses per model.
+The prompt stability was tested by rearranging the respiratory therapy categorization prompt (since it was large and complex) and evaluating whether the model returned the same answer.
 The changes made to the prompt were the following: 1) Moved the instructions after the concept criteria; 2) Reversed the ordering of the concept-related questions; and 3) Reversed the ordering of the concepts within the prompt.
@@
 The resource requirements comparison for both methods is presented in Table 4.
-From this comparison, we note that while the LLM-based methods required more specialized hardware and software, the time savings for classification was estimated to be significant when compared to manual review in the traditional approach, especially when considering that the actual concept classification is performed automatically by the LLM.
+From this comparison, we note that while the LLM-based methods required more specialized hardware and software, the time savings for classification was estimated to be significant when compared to manual review in the traditional approach, especially when considering that the actual concept classification is performed by the model.
 Therefore, we decided LLM-based methods would be appropriate to explore for concept classification.
@@
 Discussion
-In this study, we developed PHEONA, an evaluation framework specifically for LLM applications to computational phenotyping.
+In this study, we developed PHEONA, an evaluation framework specifically for applying LLMs to computational phenotyping.
 In addition to outlining components of PHEONA, we demonstrated how to apply it to concept classification, a single task within a broader phenotyping pipeline for ARF respiratory support therapies.
+As LLM-based methods continue to evolve, we encourage corresponding modifications to PHEONA to support these changes.
 Evaluation Framework
-The primary contribution of this study was development of PHEONA to support advancement of computational phenotyping methods by integrating LLMs.
-Despite recent improvements, computational phenotyping remains complicated and costly[shang_making_2019, tekumalla_towards_2024] and many of the desiderata for computational phenotyping identified over a decade ago are still relevant today, indicating the continued need for substantial improvements[hripcsak_next-generation_2013, wen_impact_2023].
+The primary contribution of this study was development of PHEONA to support advancement of computational phenotyping methods using LLMs.
+Despite recent improvements, computational phenotyping remains complicated and costly[shang_making_2019, tekumalla_towards_2024] with many of the desiderata for computational phenotyping identified over a decade ago still relevant today[hripcsak_next-generation_2013, wen_impact_2023].
 Foundation models like LLMs are the first widely available tools with the capability to comprehend and produce human-readable text for a fraction of the development time and thus, they are a natural next step for computational phenotyping.
-With PHEONA, we provided guidelines to support integration of LLMs into phenotyping pipelines.
-The resource requirements component of PHEONA was designed to assess whether LLMs are appropriate for the task of interest since not every task could or should use an LLM based on available resources.
-The model ability component was then designed to determine if the available LLMs could accomplish the task of interest.
-To simplify development and avoid additional costs, we recommend researchers start with smaller parameter LLMs and consider larger models if performance is low.
-There are other methods to consider, such as fine-tuning or multiple model approaches[li_llms-as-judges_2024], but these may not provide an advantage over traditional methods and should be evaluated using resource requirements criteria on a case-by-case basis.
-In scenarios where the resource requirements are revisited, the application of PHEONA would look similar to Figure 2, where the resource requirements and model ability are cycled through to find the optimal method for a given task.
+With PHEONA, we provided guidelines to evaluate LLM integration into phenotyping pipelines.
+The resource requirements component was designed to assess whether LLMs are appropriate for the task of interest since not every task could or should use LLMs based on available resources.
+The model ability component was then designed to determine if available LLMs could accomplish the task of interest.
+To simplify development and avoid additional costs, we suggest starting with smaller parameter LLMs and consider larger models if performance is low.
+There are other methods to consider, such as fine-tuning[li_llms-as-judges_2024], but these may not provide an advantage over traditional methods and should be evaluated on a case-by-case basis.
+In these scenarios, the application of PHEONA would look similar to Figure 2, where the resource requirements and model ability are cycled through to find the optimal method.
 Phenotyping Use Case
-Another contribution of this study was the initial results of concept classification for ARF respiratory support therapies.
-From our evaluation of the model ability criteria, we concluded that we could accomplish concept classification effectively with either Mistral or Phi, although the models we tested performed well at classifying the constructed concept sample.
-All of the tested models had high response correctness scores and thus, we only focused on secondary metrics (model response latency and response format accuracy) for final model evaluation.
-With respect to these metrics, DeepSeek performed marginally worse than the other models; however, it was also the only model to incorporate reasoning into the response generation, which makes the responses less consistent and more lengthy but potentially more valuable for complex tasks.
+Another contribution was an initial assessment of concept classification for ARF respiratory support therapies.
+From our model ability criteria evaluation, we concluded either Mistral or Phi could effectively perform concept classification.
+Since all of the models had high response correctness scores, we only focused on secondary metrics (model response latency and response format accuracy) for model evaluation.
+With respect to these metrics, DeepSeek performed marginally worse than the other models; however, it was also the only model to incorporate reasoning into its response, which makes the responses less consistent and more lengthy but potentially more valuable for complex tasks.
 These results demonstrated several advantages of LLM-based approaches specifically for classification tasks.
-First, these results showed that high accuracy can be achieved without extensive feature engineering or model training.
+First, these results showed high accuracy can be achieved without extensive feature engineering or model training.
 Second, our methods incurred no additional hardware costs, indicating that LLM-based methods do not need expensive workflows to be effective for classification.
@@
 Study Limitations
-There are some limitations to the framework and methods identified in this study.
-First, there are known issues with using LLMs, including the ability to produce hallucinations and reduced performance around specialized, domain-specific knowledge[li_llms-as-judges_2024, patil_review_2024].
+There are known issues with using LLMs, including the ability to produce hallucinations and reduced performance around specialized, domain-specific knowledge[li_llms-as-judges_2024, patil_review_2024].
 While the model ability components of PHEONA help highlight where LLMs may fail, there may be cases where LLMs cannot accomplish the task of interest and traditional phenotyping methods must be used instead.
 Additionally, the current version of PHEONA requires manual review of responses to measure each criterion, despite the known drawback of manual review being time-intensive.
-Although we used a random sample of the total constructed concepts to make manual review more feasible, it is possible that performance on the entire dataset is not representative of performance on the random sample.
+Although we used a random sample of constructed concepts to make manual review more feasible, it is possible performance on the entire dataset is not representative of performance on the random sample.
 Finally, for our use case, we only considered the binary classification task of “YES/NO” for whether each constructed concept was relevant to the topics of interest although multi-class classification may be more useful for downstream processes.
-However, when reviewing the responses to individual classification steps in our prompt, we did not find any inaccurate responses and thus, we should be able to convert from binary to multi-class classification with some minor prompt updates.
+However, since we did not observe any inaccurate responses when reviewing responses to individual classification steps within our prompt, we should be able to convert from binary to multi-class classification using prompt updates.
+For example, we could ask the model to return the most relevant therapy (IMV, NIPPV, or HFNI) rather than “YES/NO.”
 Future Directions
-There were several future directions identified from this study.
-First, since one of the limitations of this study was the reliance on manual review for evaluation, there are opportunities to develop automated LLM processes for computational phenotyping and to update PHEONA to incorporate automated evaluation processes.
+Since one of the limitations of this study was the reliance on manual review for evaluation, there are opportunities to develop automated LLM agents for computational phenotyping and update PHEONA to incorporate automated evaluation processes[qiu_llm-based_2024].
 Furthermore, while we demonstrated PHEONA on a single task, the framework can be generalized to adapt LLM-based methods to entire computational phenotyping pipelines rather than simply adapting LLMs to individual phenotyping tasks.
 Finally, each of the criteria in PHEONA can be expanded upon to include a more standardized evaluation approach.
-For example, there have been frameworks developed to measure LLM sensitivity to general inputs that can be adapted to the prompt stability component of PHEONA [zhuo_prosa_2024].
+For example, a framework developed to measure LLM sensitivity to general inputs can be adapted to the prompt stability component of PHEONA [zhuo_prosa_2024].
 Conclusions
 We developed PHEONA, an evaluation framework for applying LLM-based methods to computational phenotyping tasks.
-This framework can be used to assess whether LLMs are feasible and if so, which model is best for the task of interest.
-We provided an example use case to demonstrate the framework and promote further research into improving traditional computational phenotyping methods.
-[…]
+This framework helps assess if LLMs are feasible and if so, which model is best for each task.
+We demonstrated PHEONA on an example use case to promote further research into improving computational phenotyping.
 Table 1: Evaluation criteria for resource requirements, specifically attributes of development effort and development schedule.
@@
 May be re-evaluated to accommodate any inference optimization strategies implemented[chitty-venkata_survey_2023].
-Assess the maximum computational resources in each of the available host environments, focusing on graphical processing unit (GPU)resources in particular.
+Assess the maximum computational resources in each of the available host environments, focusing on graphics processing unit (GPU)resources in particular.
 Assess whether maximum resources will be available throughout the project or if a smaller model size is desired to ensure consistent access.
@@
 Accuracy How accurate the model is at classifying or phenotyping based on a ground truth.
+Based on task requirements, this may also include assessing the accuracy of responses with respect to the reference text or provided context.
 Classification accuracy can be measured using common classification metrics like accuracy, precision, recall, and area under the receiver operating characteristic (AUROC).
+Accuracy against reference text can be performed using N-gram analysis, BLEU, or ROUGE while accuracy against provided context can be performed also by N-gram analysis and additionally ROUGE-C.[sai_survey_2022]
 Quasi-Accuracy Classification accuracy after marking relevant or semantically similar false positives as true positives to avoid penalizing LLM when unnecessary.
@@
 Time for Manual Review An estimated $66$ months of the total $1818$ months for ARF phenotype development were required for manual review of concepts.
-An week was required for clinical experts to collate information on the concepts of interest to be used for prompt engineering and an additional day was required to manually review model ability results.
+A week was required for clinical experts to collate information on the concepts of interest to be used for prompt engineering and an additional day was required to manually review model ability results.
 LLM-based method was superior.
@@
 a AUROC: area under the receiver operating characteristic curve.
+[…]
+[heading=subbibliography]